# Travelling waves for Maxwell’s equations in nonlinear and symmetric media

### v2

@@ Introduction
defined on a Banach space $X\subset L^{2}(\mathbb{R}^{2})^{6}\cap L^{p}(\mathbb{R}^{2})^{6}$ given later, where $p>2$. Here, $b_{L}(\cdot,\cdot)$ is the bilinear form associated with $L$, i.e. $b_{L}(u,\varphi)=\int_{\mathbb{R}^{2}}\langle Lu,\varphi\rangle\,dx$ for all $u\in\operatorname{dom}(L)$ and all $\varphi\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$.
@@
Suppose that (V), (F0)–(F3) are satisfied and $F$ is radial. Then there exist infinitely many solutions to (1.7) of the form $u_{n}=v_{n}+w_{n}$ such that $J(u_{n})\to\infty$ as $n\to\infty$, $v_{n}\in H^{1}(\mathbb{R}^{2})^{6}$, $v_{n}\neq 0$, $w_{n}\in L^{2}(\mathbb{R}^{2})^{6}\cap L^{p}(\mathbb{R}^{2})^{6}$, $Lw_{n}=0$, and $g\star u_{n}=u_{n}$. Moreover, each $u_{n}=\begin{pmatrix}U_{n}\\ \widetilde{U}_{n}\end{pmatrix}$ has the profiles of the form
@@ 3. Analysis of Cerami sequences
$\Big{|}\overset{\circ}{\nabla}\times\begin{pmatrix}(g\star U)(x)\\ (g\star\widetilde{U})(x)\end{pmatrix}\Big{|}=\Big{|}\overset{\circ}{\nabla}\times\begin{pmatrix}U(g^{-1}x)\\ \widetilde{U}(g^{-1}x)\end{pmatrix}\Big{|}$
@@
First, we claim that ${\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ is compactly embedded into $L^{p}(\mathbb{R}^{2})^{6}$ for any $p>2$. Indeed, let $(v_{n})\subset{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ be such that $v_{n}\rightharpoonup 0$ as $n\to\infty$. Passing to a subsequence $v_{n}\to 0$ a.e. on $\mathbb{R}^{2}$ as $n\to\infty$. Then, $(|v_{n}|)$ is bounded in $H^{1}(\mathbb{R}^{2})$ and ${\mathcal{O}}(2)$-invariant. Hence, $|v_{n}|\to v_{0}$ in $L^{p}(\mathbb{R}^{2})$ along a subsequence as $n\to\infty$, which yields $v_{0}=0$.
@@ 4. Non-TE modes
$u_{\rho}(x)=\begin{pmatrix}\alpha(x)x_{1}\\ \alpha(x)x_{2}\\ 0\\ \widetilde{\alpha}_{1}(x)\\ \widetilde{\alpha}_{2}(x)\\ \widetilde{\alpha}_{3}(x)\end{pmatrix},\quad u_{\tau}(x)=\begin{pmatrix}-\beta(x)x_{2}\\ \beta(x)x_{1}\\ 0\\ 0\\ 0\\ 0\end{pmatrix},\quad u_{\zeta}(x)=\begin{pmatrix}0\\ 0\\ \gamma(x)\\ \widetilde{\gamma}_{1}(x)\\ \widetilde{\gamma}_{2}(x)\\ \widetilde{\gamma}_{3}(x)\end{pmatrix}$
for some $\widetilde{\alpha}_{1},\widetilde{\alpha}_{2},\widetilde{\gamma}_{2},\widetilde{\gamma}_{2}$ and radial $\alpha,\beta,\gamma,\widetilde{\alpha}_{3},\widetilde{\gamma}_{3}$. Note that
@@
$\overset{\circ}{\nabla}\times u_{\zeta}(x)=\begin{pmatrix}\partial_{x_{2}}\gamma-k\widetilde{\gamma}_{2}\\ k\widetilde{\gamma}_{1}-\partial_{x_{1}}\gamma\\ \ 0\\ \partial_{x_{2}}\widetilde{\gamma}_{3}\\ -\partial_{x_{1}}\widetilde{\gamma}_{3}\\ \partial_{x_{1}}\widetilde{\gamma}_{2}-\partial_{x_{2}}\widetilde{\gamma}_{1}\end{pmatrix}$
@@
$\langle\overset{\circ}{\nabla}\times u_{\tau}(x),\overset{\circ}{\nabla}\times u_{\zeta}(x)\rangle=0$
since $\widetilde{\gamma}_{3}$ is radial. Therefore, (4.1) holds and
@@
We check (i) first. Fix $\varepsilon\in(0,\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V)$ and let $C_{\varepsilon}>0$ such that (3.1) holds. Then, denoting $c_{\varepsilon}:=\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V-\varepsilon>0$, from (3.2), for every $v\in{\mathcal{V}}$ there holds
@@
Let $\delta\in(0,c_{\varepsilon}/C_{\varepsilon})$. From (N2), there exists $r>0$ such that $\Phi(t)\leq\delta t^{2}$ for all $|t|\leq r$. Now, let $\kappa>1$ be the constant appearing in the characterization of the $\Delta_{2}$ condition, that is, $t\Phi^{\prime}(t)\leq\kappa\Phi(t)$ for all $t\in\mathbb{R}$, cf. Lemma 2.1. Then, there exists $K>0$ such that $\Phi(t)\leq K|t|^{\kappa}$ for all $|t|\geq\delta$. Observe that this and (N3) yield $\kappa>2$. There holds
@@
$\beta_{n}:=\sup\left\{|v|_{p}:v\in Z_{n}\text{ and }\|v\|=1\right\}$

### v3

@@ Introduction
defined on a Banach space $X\subset L^{2}(\mathbb{R}^{2})^{6}$ given later, where $p>2$. Here, $b_{L}(\cdot,\cdot)$ is the bilinear form associated with $L$, i.e. $b_{L}(u,\varphi)=\int_{\mathbb{R}^{2}}\langle Lu,\varphi\rangle\,dx$ for all $u\in\operatorname{dom}(L)$ and all $\varphi\in{\mathcal{C}}_{0}^{\infty}(\mathbb{R}^{2})^{6}$.
@@
Suppose that (V), (F0)–(F3) are satisfied and $F$ is radial. Then there exist infinitely many solutions to (1.7) of the form $u_{n}=v_{n}+w_{n}$ such that $J(u_{n})\to\infty$ as $n\to\infty$, $v_{n}\in H^{1}(\mathbb{R}^{2})^{6}$, $v_{n}\neq 0$, $w_{n}\in L^{2}(\mathbb{R}^{2})^{6}$, $\Phi(w_{n})\in L^{1}(\mathbb{R}^{2})$, $Lw_{n}=0$, and $g\star u_{n}=u_{n}$. Moreover, each $u_{n}=\begin{pmatrix}U_{n}\\ \widetilde{U}_{n}\end{pmatrix}$ has the profiles of the form
@@ 3. Analysis of Cerami sequences
[…]
@@
First, we claim that ${\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ is compactly embedded into $L^{r}(\mathbb{R}^{2})^{6}$ for any $r>2$. Indeed, let $(v_{n})\subset{\mathcal{V}}_{{\mathcal{S}}{\mathcal{O}}(2)}$ be such that $v_{n}\rightharpoonup 0$ as $n\to\infty$. Passing to a subsequence $v_{n}\to 0$ a.e. on $\mathbb{R}^{2}$ as $n\to\infty$. Then, $(|v_{n}|)$ is bounded in $H^{1}(\mathbb{R}^{2})$ and ${\mathcal{O}}(2)$-invariant. Hence, $|v_{n}|\to v_{0}$ in $L^{r}(\mathbb{R}^{2})$ along a subsequence as $n\to\infty$, which yields $v_{0}=0$.
@@ 4. Non-TE modes
$u_{\rho}(x)=\begin{pmatrix}\alpha(x)x_{1}\\ \alpha(x)x_{2}\\ 0\\ \widetilde{\alpha}_{1}(x)\\ \widetilde{\alpha}_{2}(x)\\ \widetilde{\alpha}_{3}(x)\end{pmatrix},\quad u_{\tau}(x)=\begin{pmatrix}-\beta(x)x_{2}\\ \beta(x)x_{1}\\ 0\\ 0\\ 0\\ 0\end{pmatrix},\quad u_{\zeta}(x)=\begin{pmatrix}0\\ 0\\ \gamma(x)\\ 0\\ 0\\ 0\end{pmatrix}$
for some $\widetilde{\alpha}_{1},\widetilde{\alpha}_{2}$, and radial $\alpha,\beta,\gamma,\widetilde{\alpha}_{3}$. Note that
@@
$\overset{\circ}{\nabla}\times u_{\zeta}(x)=\begin{pmatrix}\partial_{x_{2}}\gamma\\ -\partial_{x_{1}}\gamma\\ 0\\ 0\\ 0\\ 0\end{pmatrix}$
@@
$\langle\overset{\circ}{\nabla}\times u_{\tau}(x),\overset{\circ}{\nabla}\times u_{\zeta}(x)\rangle=0.$
Therefore, (4.1) holds and
@@
We check (i) first. Fix $\varepsilon\in\left(0,\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V\right)$ and let $C_{\varepsilon}>0$ such that (3.1) holds. Then, denoting $c_{\varepsilon}:=\frac{1}{2}k^{2}-\frac{1}{2}\operatorname*{ess\,sup}V-\varepsilon>0$, from (3.2), for every $v\in{\mathcal{V}}$ there holds
@@
Let $\delta\in(0,c_{\varepsilon}/C_{\varepsilon})$. From (N2), there exists $r>0$ such that $\Phi(t)\leq\delta t^{2}$ for all $|t|\leq r$. Now, let $\kappa>1$ be the constant appearing in the characterization of the $\Delta_{2}$ condition, that is, $t\Phi^{\prime}(t)\leq\kappa\Phi(t)$ for all $t\in\mathbb{R}$, cf. Lemma 2.1. Then, there exists $K>0$ such that $\Phi(t)\leq K|t|^{\kappa}$ for all $|t|\geq r$. Observe that this and (N3) yield $\kappa>2$. There holds
@@
$\beta_{n}:=\sup\left\{|v|_{\kappa}:v\in Z_{n}\text{ and }\|v\|=1\right\}$